Five times a number less the square of the number is equal to the negative of the number. Determine the number.
Answer by swincher4391(1107)

Let n be the number.
5n - n^2 = -n
n^2 - 6n = 0
n(n-6) = 0
n = 0
n = 6

Check 0 - 0 = - (0)
5*6 - 6^2 = -6